Let The National- Income Model Be: Y = C + I0 + G) C= A+B +B(Y- T0) (A ˃ 0, 0 ˂ B ˂ 1) G = Gy (0 ˂ G ˂ 1) (A) Identify The Endogenous Variables. (B) Give The Economic Meaning Of The Parameter G. (C) Find The Equilibrium National Income. (D) What Restriction On The


This problem has been solved!

See the answer

5. Let the national- income model be:

Y = C + I0 + G)

C= a+b +b(Y- T0) (a ˃ 0, 0 ˂ b ˂ 1)

G = gy (0 ˂ g ˂ 1)

(a) Identify the endogenous variables.

(b) Give the economic meaning of the parameter g.

(c) Find the equilibrium national income.

(d) What restriction on the parameters is needed for a solution to exits?



1
Expert's answer
2021-10-17T17:49:04-0400

(a) The endogenous variables are:

Y, C. Values for these variables are only determined by solving the model.

(b) The parameter g is for marginal propensity of government expenditure.

(c) The equilibrium national income is:

Y = a + b + b(Y - T0) + I0 + gY,

Y(1 - b - g) = a + b - bT0 + I0,

"Y =\frac{a + b - bT0 + I0}{1 - b - g}."

(d) The restriction on the parameters ineeded for a solution to exits are:

"1 - b - g \neq 0" or "b + g \neq 1" .